OpenStudy (anonymous):

Using complete sentences, explain how to find the zeros of the function f(x) = 2x^3 – 9x + 3.

OpenStudy (radar):

Type in the URL of Wolfram Alpha Type the function the box at top of page and click on the = sign.

OpenStudy (radar):

There will be three roots. I have no idea how to isolate them when the function is prime.

OpenStudy (campbell_st):

1st using the remainder theorem find P(0) = 3 next use the remainder theorem again to find P(1) = -4 so a root exists between x = 0 and x = 1 use Newton's method to approximate the roots of a polynomial \[x _{1} = x _{0} - f(x _{0})/f'(x _{0})\] let \[x _{0} = 1/2\] f'(x) = 6x^2 - 9x f(1/2) = \[x _{1}\approx 1/2 - (-1.25)/(-7.5)\] \[x _{1}\approx 1/3\] another root exists between -2 and -3 P(-2) = 5 P(-3) = -24 axxproximate let x = -2.5 and then use newton's method

OpenStudy (radar):

The three roots as found by Wolfram are: -2.27163, 0.342241, and 1.92939

OpenStudy (campbell_st):

a root also exists between P(1) = -4 and P(2) = 1 find an appropriate approximation then use Newton's method again